OpenStudy (anonymous):

The point (-1,2) lies on a circle whose center is at (-4,1) . Find an equation of this circle.

OpenStudy (anonymous):

first find the distance. that is the radius of the circle. distance is \[\sqrt{(-1+4)^2+(2-1)^2}\] \[\sqrt{3^2+1^2}=\sqrt{10}\] then just write \[(x+4)^2+(y-1)^2=10\]